### v2

# A probabilistic proof of Euler’s pentagonal number theorem

Abstract.
In this paper, we present a probabilistic proof of Euler’s pentagonal number theorem based on a shuffling model.
Key words and phrases:
Euler’s pentagonal number theorem, probabilistic proof, shuffling model.
2020 Mathematics Subject Classification:
60C05, 05A19, 11P84.
1. Prologue
Euler’s pentagonal number theorem is recognized as one of the pioneering discoveries in the analysis of the infinite. Formally, it is the following series expansion:
Theorem 1.1.
$\prod_{j\geq 1}(1-q^{j})=\sum_{n=-\infty}^{\infty}(-1)^{n}q^{\frac{n(3n+1)}{2}}.$ (1.1)
This result was first communicated from Euler to Daniel Bernoulli in a letter dated November 30, 1740111Euler’s original letter, unfortunately, has not been preserved but the lines regarding the pentagonal number theorem were alluded in Bernoulli’s reply dated January 28, 1741. See [5, Brief 52].; however, no proof was included therein. The seemingly earliest demonstration offered by Euler [3] was published in 1760, which is a retelling of the idea in his letter to Goldbach [4, Brief 144] on June 9, 1750. Euler’s proof, as briefly presented at the end of this paper, is inductive, relying on an ingenious arithmetic manipulation of the series.
Ever since its discovery, the pentagonal number theorem has shown broad and profound connections to other fields. As observed by Legendre [8, pp. 132–133, §458], the series expansion (1.1) may be combinatorially interpreted in terms of integer partitions; this interpretation motivated Franklin’s bijective proof [6] half a century later. Also, Goncharova [7] noticed that the pentagonal number theorem arises as a natural implication from the calculation of the cohomology of the negative part of the Witt algebra, and it is just a tip of a big picture that bonds Lie algebra, partition theory, and $q$-series. We refer the interested reader to a terse exposition by Andrews [1], and a more detailed summary on Euler’s discovery by Bell [2].
The objective of the present paper is to understand the pentagonal number theorem in a different way. To be specific, we propose a shuffling model as given in the next section and utilize it to provide a new probabilistic proof.
2. A shuffling model
Let $\mathbb{N}_{+}$ be the set of positive integers, and :=+$\mathbb{N}:=\mathbb{N}_{+}\cup\{0\}$ the set of natural numbers. Given $a,b\in\mathbb{N}$, we denote by $[a,b]$ the set of natural numbers $n$ such that $a\leq n\leq b$, and by $[a,\infty)$ the set of natural numbers $n$ such that $n\geq a$.
Throughout, we always assume that $q$ is a real number such that $10<q<$ 1.
Recall that for $S$ a subset of $\mathbb{N}$, we say a sequence $\boldsymbol{s}:=(s_{0},s_{1},s_{2},\ldots)$ is a shuffle of $S$ if every number in $\boldsymbol{s}$ belongs to $S$ and every number in $S$ appears exactly once in $\boldsymbol{s}$.
Now given a sequence $(s_{0},\ldots,s_{k-1},s_{k})$ of $k+1$ different natural numbers for a certain $k\geq 1$, we define a probability model k as follows:
k $:(s_{0},\ldots,s_{k-1},s_{k})$
$\ \ \mapsto\begin{cases}(s_{0},\ldots,s_{k-1},s_{k})&\text{probability $1-q$},\\ (s_{0},\ldots s_{i-1},s_{k},s_{i},\ldots,s_{k-1})&\text{probability $q^{k-i}(1-q)$ for $i\in[1,k-1]$},\\ (s_{k},s_{0},\ldots,s_{k-1})&\text{probability $q^{k}$}.\end{cases}$
Note that the outcomes are shuffles of the original sequence. Intuitively, we swap $s_{k}$ with its preceding entry with a probability $q$ through a step-by-step procedure, and the program is terminated when a certain step of swapping is not executed or $s_{k}$ already reaches the leftmost position.
We then describe our shuffling model :
  • Round $0$: We begin with the sequence $(0)$;
  • Round $k$ ($k\geq 1$): For the resulting sequence of the previous round, we append the number $k$ to the right end (so that the new sequence is of length $k+1$) and then apply the probability model k.
It is notable that the sample space of our model covers all shuffles of $\mathbb{N}$. For example, the sequence $(2,0,3,1,4,5,6,7\ldots)$ is obtained by making no swap in Round $1$, two swaps in Round $2$, one swap in Round $3$, and no swap in all further rounds.
3. Event spaces
To provide a clear description of the event spaces of our interest, it is necessary to categorize the event space of each round. Recalling the definition of the model k, we label the sequences in the sample space as follows:
$\boldsymbol{s}_{i}:=\begin{cases}(s_{0},\ldots,s_{k-1},s_{k})&\text{for $i=0$},\\ (s_{0},\ldots s_{k-i-1},s_{k},s_{k-i},\ldots,s_{k-1})&\text{for $i\in[1,k-1]$},\\ (s_{k},s_{0},\ldots,s_{k-1})&\text{for $i=k$}.\end{cases}$
We say
  • The shuffling k is leftmost if the event space is $\{\boldsymbol{s}_{k}\}$;
  • The shuffling k is anti-leftmost if the event space is $\{\boldsymbol{s}_{0},\ldots,\boldsymbol{s}_{k-1}\}$;
  • The shuffling k is free if the event space is $\{\boldsymbol{s}_{0},\ldots,\boldsymbol{s}_{k-1},\boldsymbol{s}_{k}\}$.
Now for our shuffling model , we are only interested in the event spaces wherein each round belongs to one of the three types. To begin with, we assign a label for the $k$-th round:
  • $k_{\triangleright}$ means k is leftmost;
  • $k_{\triangleleft}$ means k is anti-leftmost;
  • $k_{\lozenge}$ means k is free.
Our desired event space is a collection of shuffles denoted by the set
$\big{\langle}1_{\square_{1}},2_{\square_{2}},3_{\square_{3}},\ldots\big{\rangle},$ (3.1)
wherein each $\square_{k}\in\{\triangleright,\triangleleft,\lozenge\}$, meaning that in the $k$-th round the shuffling k is of type $k_{\square_{k}}$.
We may further collect the rounds that are leftmost and anti-leftmost, respectively, and the two sets of rounds are written as $I$ and $J$. It is clear that $I$ and $J$, as subsets of $\mathbb{N}_{+}$, are disjoint. Also, the set of free rounds is given by the complement of the disjoint union $I\sqcup J$ with respect to $\mathbb{N}_{+}$. Hence the event space (3.1) may be abbreviated as
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}.$ (3.2)
We illustrate the above notation with an example. Consider the sequence
$(6,1,0,4,2,3,5,7,8,9,10,\ldots).$
It belongs to the event spaces $\big{\langle}\varnothing_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$, $\big{\langle}\{1\}_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$, $\big{\langle}\{1,6\}_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$, ,$\big{\langle}\varnothing_{\triangleright},\{2\}_{\triangleleft}\big{\rangle}$, ,$\big{\langle}\varnothing_{\triangleright},\{4,5\}_{\triangleleft}\big{\rangle}$, $\big{\langle}\{1\}_{\triangleright},\{2,4\}_{\triangleleft}\big{\rangle}$ (and many others) but not the space $\big{\langle}\{2\}_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$ (because in this space $2$ should appear at the leftmost position in the subsequence projected onto $\{0,1,2\}$) nor ,$\big{\langle}\varnothing_{\triangleright},\{1\}_{\triangleleft}\big{\rangle}$ (because in this space $1$ should not appear at the leftmost position in the subsequence projected onto $\{0,1\}$).
We close this section with two observations.
Proposition 3.1.
Let $I$ and $J$ be two disjoint subsets of $\mathbb{N}_{+}$ and $I^{\prime}$ and $J^{\prime}$ another two disjoint subsets of $\mathbb{N}_{+}$. If there exists a positive integer $k$ such that $k\in I$ and $k\in J^{\prime}$, then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}\cap\big{\langle}I^{\prime}_{\triangleright},J^{\prime}_{\triangleleft}\big{\rangle}=\varnothing.$ (3.3)
Proof.
Let $\boldsymbol{s}\in\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}$ and $\boldsymbol{s}^{\prime}\in\big{\langle}I^{\prime}_{\triangleright},J^{\prime}_{\triangleleft}\big{\rangle}$ be arbitrary. We project $\boldsymbol{s}$ onto $[0,k]$. Since the shuffling k with respect to $\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}$ is leftmost, the number $k$ must appear at the leftmost position in the projection. Meanwhile, in the projection of $\boldsymbol{s}^{\prime}$ onto $[0,k]$, the number $k$ cannot appear at the leftmost position because the shuffling k with respect to $\big{\langle}I^{\prime}_{\triangleright},J^{\prime}_{\triangleleft}\big{\rangle}$ is anti-leftmost. Hence, $\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}$ and $\big{\langle}I^{\prime}_{\triangleright},J^{\prime}_{\triangleleft}\big{\rangle}$ are disjoint. ∎
Next, we recall that the union of two sets $S_{1}$ and $S_{2}$ is written as $S_{1}\sqcup S_{2}$ once $S_{1}$ and $S_{2}$ are disjoint; this union is called a disjoint union.
Proposition 3.2.
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$. Then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}=\bigsqcup_{K^{\prime}\subset K}\big{\langle}(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\big{\rangle}.$ (3.4)
In particular, if $K$ is a singleton, namely, $K=\{k\}$ for a certain $k\in\mathbb{N}_{+}$, then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}=\big{\langle}(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\big{\rangle}\sqcup\big{\langle}I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\big{\rangle}.$ (3.5)
Proof.
We only need the fact that for each $k\in K$, the event space of a free shuffling k is the disjoint union of the event spaces of a leftmost k and an anti-leftmost k. ∎
4. Probabilities
While we have determined the event spaces for our investigation, the next task is to analyze the probabilities. Throughout, for a probability model, we denote by $\mathbf{Pr}_{E}$ the total probability of all outcomes in the event space $E$.
Recalling again the definition of the model k, we have
[…]
The following evaluation is then immediate.
Proposition 4.1.
Let $I$ and $J$ be disjoint subsets of $\mathbb{N}_{+}$. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\prod_{i\in I}q^{i}\prod_{j\in J}(1-q^{j}).$ (4.1)
Proposition 4.1 has two implications.
Corollary 4.2.
Let $I$ and $J$ be two disjoint subsets of $\mathbb{N}_{+}$ with $I$ being infinite. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=0.$ (4.2)
Corollary 4.3.
Let $I$ and $I^{\prime}$ be two finite subsets of $\mathbb{N}_{+}$, not necessarily disjoint, and let $J$ be a subset of $\mathbb{N}_{+}$ that is disjoint with both $I$ and $I^{\prime}$. Suppose that
$\sum_{i\in I}i=\sum_{i^{\prime}\in I^{\prime}}i^{\prime}.$
Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle I^{\prime}_{\triangleright},J_{\triangleleft}\rangle}.$ (4.3)
Finally, we recall that in a probability model, if two event spaces $E_{1}$ and $E_{2}$ are disjoint, then
$\mathbf{Pr}_{E_{1}\sqcup E_{2}}=\mathbf{Pr}_{E_{1}}+\mathbf{Pr}_{E_{2}}.$
Therefore, Proposition 3.2 implies the following relation.
Proposition 4.4.
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\sum_{K^{\prime}\subset K}\mathbf{Pr}_{\langle(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\rangle}.$ (4.4)
In particular, if $K$ is a singleton, namely, $K=\{k\}$ for a certain $k\in\mathbb{N}_{+}$, then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\rangle}+\mathbf{Pr}_{\langle I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\rangle}.$ (4.5)
5. The pentagonal number theorem
The key to our proof of the pentagonal number theorem is the following relation.
Theorem 5.1.
For every nonnegative integer $N$,
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\sum_{n=0}^{N}(-1)^{n}\big{(}\mathbf{Pr}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}$
$\quad-(-1)^{N}\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}.$ (5.1)
We show this result by induction on $N$.
Base case.
We need to prove the case where $N=0$:
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle\{1\}_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{Pr}_{\langle\{k+1\}_{\triangleright},[1,k]_{\triangleleft}\rangle}$
$=\mathbf{Pr}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{Pr}_{\langle\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\rangle}.$
Note that $\big{\langle}\varnothing_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$ contains all shuffles of $\mathbb{N}$. Meanwhile, given an arbitrary shuffle, either none of the rounds are leftmost (which is characterized by $\big{\langle}\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\big{\rangle}$) or there exists a unique round $k\geq 1$ such that k is the first leftmost shuffling (which is characterized by $\big{\langle}\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\big{\rangle}$). Furthermore, given two different positive integers $k$ and $k^{\prime}$, we assume without loss of generality that $k<k^{\prime}$. Then $k\in\{k\}$ and $k\in[1,k^{\prime}-1]$. In light of Proposition 3.1,
$\big{\langle}\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\big{\rangle}\cap\big{\langle}\{k^{\prime}\}_{\triangleright},[1,k^{\prime}-1]_{\triangleleft}\big{\rangle}=\varnothing.$
Hence,
$\big{\langle}\varnothing_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}=\big{\langle}\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\big{\rangle}\sqcup\left(\bigsqcup_{k\geq 1}\big{\langle}\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\big{\rangle}\right).$
Applying Proposition 4.4 leads us to the required relation. ∎
Inductive step.
It suffices to show that for each nonnegative integer $N$,
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+3+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
We start by noting that for each $k\geq 1$, the sets $[k+1+N,k+1+2N]$, $[2+N,k+N]$ and $\{1+N\}$ are pairwise disjoint. It follows from Proposition 4.4 that
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
For the first sum on the right-hand side,
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad+\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
We further note that the sets $[k+2+N,k+2+2N]$, $[2+N,k+N]$ and $\{k+1+N\}$ are pairwise disjoint. Another application of Proposition 4.4 yields
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
Finally, we see from Corollary 4.3 that
$\sum_{k\geq 1}\mathbf{Pr}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
This implies that
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
Singling out the $k=1$ term of the sum in the above gives the desired part $\mathbf{Pr}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$. ∎
Now we are ready to conclude the pentagonal number theorem.
Proof of Theorem 1.1.
Letting $N\to\infty$ in (5.1) and recalling Corollary 4.2, we have
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}=\sum_{n\geq 0}(-1)^{n}\big{(}\mathbf{Pr}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}.$
Applying Proposition 4.1 yields
$\prod_{j\geq 1}(1-q^{j})=\sum_{n\geq 0}(-1)^{n}\big{(}q^{\frac{n(3n+1)}{2}}-q^{\frac{(n+1)(3n+2)}{2}}\big{)}.$
This is exactly (1.1). ∎
6. Epilogue
Before putting down the curtain, let us flash back to Euler’s original proof of the pentagonal number theorem in [3]. Adopting the $q$-Pochhammer symbol for $n\in\mathbb{N}\cup\{\infty\}$:
$(A;q)_{n}$ $:=\prod_{j=0}^{n-1}(1-Aq^{j}),$
the key observation of Euler, although not explicitly stated, is as follows:
Theorem 6.1.
For every nonnegative integer $N$,
$(q;q)_{\infty}$ $=\sum_{n=0}^{N}(-1)^{n}\big{(}q^{\frac{n(3n+1)}{2}}-q^{\frac{(n+1)(3n+2)}{2}}\big{)}$
$\quad-(-1)^{N}\sum_{k\geq 1}q^{\frac{(1+N)(2k+2+3N)}{2}}(q^{1+N};q)_{k}.$ (6.1)
It is worth noting that our proof can be understood as a translation of Euler’s proof into the language of classical probability theory. To tie the knot, we simply apply Proposition 4.1 and derive that
$\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\prod_{i=k+1+N}^{k+1+2N}q^{i}\prod_{j=1+N}^{k+N}(1-q^{j})$
$=q^{\frac{(1+N)(2k+2+3N)}{2}}(q^{1+N};q)_{k}.$
In view of (5.1), this is clearly the last piece of expressions required for (6.1).
Acknowledgements
This work was supported by the Austrian Science Fund (No. 10.55776/F1002). I am grateful to Hugh Thomas for the comments on Proposition 4.4 in an earlier draft.